# Efficient Time Series Processing for Transformers and State-Space Models through Token Merging

### v1

@@ Abstract
Transformer architectures have shown promising results in time series processing. However, despite recent advances in subquadratic attention mechanisms or state-space models, processing very long sequences still imposes significant computational requirements. Token merging, which involves replacing multiple tokens with a single one calculated as their linear combination, has shown to considerably improve the throughput of vision transformer architectures while maintaining accuracy. In this work, we go beyond computer vision and perform the first investigations of token merging in time series analysis on both time series transformers and state-space models. To effectively scale token merging to long sequences, we introduce local merging, a domain-specific token merging algorithm that selectively combines tokens within a local neighborhood, adjusting the computational complexity from linear to quadratic based on the neighborhood size. Our comprehensive empirical evaluation demonstrates that token merging offers substantial computational benefits with minimal impact on accuracy across various models and datasets. On the recently proposed Chronos foundation model, we achieve accelerations up to with only minor accuracy degradations.
@@ 1 Introduction
Since their inception in NLP [1], transformers have extended their influence into various domains, including computer vision with Vision Transformers (ViTs) [2], graphs [3, 4], and time series processing [5]. However, the computational complexity of the standard attention mechanism used in transformer architectures scales quadratically with the number of input tokens, resulting in high memory requirements. This scalability issue becomes especially pronounced in time series processing, where sequences frequently comprise thousands of tokens [6]. Consequently, recent foundational models in time series, such as Chronos, exhibit impressive zero-shot generalization capabilities but demand substantial computational resources [7].
@@
Recently, state-space models have emerged as a solution to mitigate the computational burden of transformers. Their complexity scales subquadratically with the sequence length [8], which allows them to process millions of tokens [9]. However, even in state-space models, very long sequences will impose considerable memory and computational demands.
@@
Bolya et al. [10], have shown that the efficiency of ViTs can be substantially improved by merging tokens throughout the transformer architecture. Specifically, they compute similarity scores between tokens and combine them into single tokens through a convex combination. However, they only explore token merging for ViT architectures.
@@
In this empirical study, we for the first time explore token merging within the time series domain. We introduce a novel local token merging algorithm whose computational complexity varies from linear to quadratic, based on the neighborhood considered for each token merge. This allows token merging to scale to long sequences. The algorithm is illustrated in figure 1. Through comprehensive empirical evaluations, we analyze the impact of token merging on various time series transformer models and state-space models. Our key contributions are as follows:
Token merging in time series We present first studies on token merging in time series analysis, exploring its application beyond transformer architectures to include state-space models. For this purpose, we propose a domain-specific token merging algorithm that combines tokens within a local neighborhood around each token. Adjusting the size of the neighborhood allows the complexity of the algorithm to range from linear to quadratic. We show that this adaptability improves upon classical token merging for long-sequence modeling tasks.
Model acceleration Our findings from both pretrained models and those trained with token merging reveal substantial computational savings with only slight reductions in accuracy across five time series transformer architectures and datasets. We also assess five different model sizes per architecture, noting greater relative accelerations in larger models.
Token merging in foundation models Specifically, token merging enhances the throughput of the time series foundation model Chronos by up to , with a marginal quality drop of in relative MSE. Across four out of five datasets, we identify Pareto optimal points where token merging simultaneously boosts throughput and increases accuracy. For example, on the ETTh1 dataset, token merging achieves a acceleration while the MSE improves by .
@@
Token merging patterns Our analysis of token merging’s effects identifies three distinct patterns: 1) a consistent decline in performance with an increasing number of merged tokens, 2) initial improvements in accuracy with few merged tokens followed by a drop as merging increases, and 3) scenarios where accuracy remains unchanged regardless of the token merging rate. We investigate all patterns and find different explanations for all behaviors.
@@ 2 Related work
Time series transformers In recent years, many transformer architectures with inductive biases for time series have been proposed, successfully outperforming classical and other deep-learning-based methods in time series forecasting quality like recurrent neural networks [5]. Most of them focus on reducing complexity by modifying the attention mechanism. LogTrans uses LogSparse attention [5] while Informer focuses only on the most relevant queries using ProbSparse attention [11]. Additionally, many architectures adopt decomposition techniques to model trend and seasonal patterns [12, 13, 14, 15]. Autoformer leverages autocorrelation as a sequence-based similarity measure in the attention mechanism [13]. FEDformer uses the frequency domain to model time series effectively [14]. Non-stationary Transformers further mitigate the effect of the time series distribution changing over time [15]. PatchTST embeds subsequences as tokens to capture local semantic information and reduce complexity [16]. Other works apply hierarchical attention [17, 18] or leverage attention between the time series variates to better model multivariate patterns [19, 20].
@@
Due to their success in the vision and NLP domain, transformer-based foundation models have recently emerged for time series, often used in zero-shot settings. Many works focus on training transformers directly on large and diverse time series datasets, usually with billions of tokens [21, 22, 23, 24]. Inspired by the success of foundation models in NLP, the recently proposed Chronos model converts continuous time series data into a fixed vocabulary and is trained on both real-world and synthetic data [7]. Besides, other research branches focus on fine-tuning vision or NLP models for time series [25] and on applying large language models directly on time series data [26].
@@
State-space models Due to the quadratic scaling of the attention mechanism, transformer architectures suffer from significant computational cost when processing very long input sequences. Recently, state-space models have shown promising results in overcoming the quadratic complexity of transformers with respect to input length. Linear state-space layers solve the sequential processing requirement of RNNs through linear state-space representations [27]. The S4 model reduces memory requirements by conditioning the state-space matrix with a low-rank correction [28]. By using implicit convolutions and a data-aware gating mechanism, Hyena [8] became one of the first state-space model architectures to match transformers on NLP tasks. Later work uses hardware-aware algorithms to improve the performance of state-space models on modern accelerators [29].
@@
Reducing tokens Many works reduce the number of processed tokens to increase the efficiency of transformer architectures in computer vision and NLP, often by pruning [30, 31]. Marin et al. [32] merge tokens in ViT architectures to reduce the loss of information associated with pruning. Bolya et al. [10] enhance the token merging algorithm, which they successfully apply to already trained models. Besides initial work on classification tasks [10], subsequent work applies token merging to diffusion models [33]. Kim et al. [34] combine merging and pruning, while other work investigates optimal merging and pruning rates [35, 36].
@@ 3 Token merging
Despite recent advances in efficient transformers, processing long input sequences still induces considerable memory requirements and computational effort. In the following, we first outline the token merging algorithm introduced by Bolya et al. [10] for ViT architectures. We then propose local merging enabling efficient token merging for state-space models and long sequence processing.
(Global) Token merging in computer vision
@@
Let a neural network consist of layers denoted as , where each layer takes the output of the previous layer as input. We assume that the input consists of tokens with dimension . Thereby, the input tokens are generated by a tokenizer out of -dimensional input data . In the computer vision domain, usually takes the form , where are the width, height, and channels of the input image, respectively, and .
@@
To improve the computational efficiency of a given model, Bolya et al. [10] combine the most similar tokens in each layer, reducing the tokens to be processed in layer to . Therefore, they split the set of all tokens into two disjoint subsets in alternation to avoid merging conflicts and allow for a parallelized computation of merging correspondences. Here and contain elements each, denoted as and respectively. The authors compute the cosine similarity between all tokens in both subsets and merge the most similar correspondences by averaging the tokens accordingly. This results in a global token merging algorithm with quadratic complexity. Lastly, the authors use a fixed to enable batch processing without needing to pad individual batch elements to the same shape after token reduction.
(Local) Token merging for time series
@@
In this work, we design token merging mechanisms for time series architectures and demonstrate run-time and even performance improvements over various datasets and models. We assume that the input consists of time stamps with variates.
@@
Previous work on token merging in image processing explored global merging schemes, where every token of each subset and could be merged with each other [10, 33].
@@
However, computing the similarity between both sets of tokens has a complexity of , which is suboptimal for sequential data often consisting of long token sequences [6, 37], and state-space models featuring subquadratic complexity [8, 9].
@@
Therefore, we propose local merging - a superset of token merging - by introducing as a locality constraint where we compute the similarity only on a local subset of tokens . Figure 1 illustrates the proposed merging algorithm. The locality constraint reduces the complexity to . Varying the locality, we achieve linear complexity by considering only neighboring tokens for merging up to quadratic complexity by considering a global merging pool, possibly exploiting more redundancy.
@@
For efficient computation, we refactor into a rectangular tensor. An upper bound for the resulting speed up can be given by . The acceleration of deeper models is expected to increase as more subsequent layers can profit from already merged tokens.
@@
Some time series transformers use processing mechanisms that require a minimum number of tokens in the forward pass. To universally enable token merging in these architectures, we further introduce as the minimum number of remaining tokens. When encountering odd numbers of tokens we exclude the most recent token from merging as we expect it to contain the most relevant information.
We derive the complexity of the token merging procedures in section A.1.
@@ 4 Experiments
We systematically explore token merging in diverse settings on time series datasets and model architectures in different sizes each. Additionally, we investigate token merging in large foundation models using Chronos in a zero-shot setting [7]. Finally, we demonstrate that token merging can be applied to state-space models for long sequence processing by using a novel local merging algorithm featuring subquadratic complexity.
Datasets We use time series forecasting datasets including ETTh1, ETTm1, Weather, Electricity and Traffic for our transformer experiments. For state-space models, we use the long-range Dummy Mouse Enhancers Ensembl dataset. See section A.2 for more details.
Model architectures For our main experiments, we use architectures, including Autoformer, FEDformer, Informer, Non-stationary Transformer, and the vanilla Transformer [1] as reference. For each model, we evaluate token merging for different model sizes with encoder layers, which we train doing hyperparameter optimization (see section A.2). We use an input length of and prediction horizon samples.
For experiments on Chronos, we use the default input length of and prediction horizon [7]. We compute the median from Chronos probabilistic forecasts and report the MSE.
For our experiments on state-space models, we use HyenaDNA medium, a genomic foundation model [9] based on the Hyena architecture [8]. We use a large input length of nucleotides utilizing Hyenas subquadratic complexity.
Applying token merging We apply token merging between the attention and MLP block in all transformer architectures as in [10]. Additionally, we only merge in the encoder and not in the decoder, as many architectures do not allow for a unified merging scheme. In architectures utilizing additional tensors like attention masks or positional biases, we merge them using the same correspondences. Many transformers exhibit quadratic attention, imposing considerable computational cost. As a result, we do not find the token merging algorithm to introduce a substantially additional overhead. Thus, we choose to profit from a global merging pool for transformer models. In state-space models, we merge tokens after the Hyena operator and choose to not introduce an operation with quadratic complexity into the architecture.
Reproducibility of measurements We report all results on the same Nvidia A6000 GPU. For training, we utilize Nvidia V100 and A100 GPUs (see section A.2). We measure the end-to-end execution time of the models using warm-ups and measurement runs per batch. The standard deviation of the execution time is generally in our experiments. Besides the execution time as practically most relevant quantity, we report FLOPs as a more hardware independent measure using the thop library [38]. We always choose the maximum possible batch size and standardize the results.
@@ 5 Results
We first present our main results for token merging on pretrained models and models trained with token merging and subsequently conduct several ablation studies.
5.1 Token merging in pretrained models
We investigate token merging on diverse time series transformer models with different inductive biases. All models are specifically trained on the target dataset and token merging is applied only during inference time, as accelerating already trained models is of high practical relevance. We choose token merging hyperparameters as described in section A.2, selecting the fastest token merging trial on the validation set that is within a increase in MSE compared to the reference without token merging.
If we do not find a trial with token merging satisfying these tight criteria, we report results without token merging, mimicking how token merging might be applied in practice. We perform all selections on the validation set and report all results on the test set.

### v3

@@ Abstract
Despite recent advances in subquadratic attention mechanisms or state-space models, processing long token sequences still imposes significant computational requirements. Token merging has emerged as a solution to increase computational efficiency in computer vision architectures. In this work, we perform the first investigations of token merging in time series analysis on both transformers and state-space models. We further introduce local merging, a domain-specific token merging algorithm that selectively combines tokens within a local neighborhood, achieving two major benefits: a) Local merging can adjust its computational complexity from quadratic to linear based on the neighborhood size to effectively scale to long sequences; b) Local merging is the first causal merging scheme enabling token merging in transformer decoders. Further, we identify spectral properties of the input data that reliably predict the potential benefits of local merging without requiring evaluation on downstream tasks. Our comprehensive empirical evaluation demonstrates that local merging offers substantial efficiency gains with minimal impact on accuracy, achieving up to acceleration on the recently proposed Chronos foundation model.
@@ 1 Introduction
Since their inception in NLP (Vaswani et al., 2017), transformers have extended their influence into various domains, including computer vision with Vision Transformers (ViTs) (Dosovitskiy et al., 2021), graphs (Yun et al., 2019), and time series processing (Li et al., 2019). However, the computational complexity of the standard attention mechanism scales quadratically with the number of input tokens, resulting in high memory requirements. This scalability issue becomes especially pronounced in time series processing, where sequences frequently comprise thousands of tokens (Godahewa et al., 2021). Consequently, recent foundational models in time series, such as Chronos, exhibit impressive zero-shot generalization capabilities but demand substantial computational resources (Ansari et al., 2024).
@@
Recently, state-space models have emerged as a solution to mitigate the computational burden of transformers. Their complexity scales subquadratically with the sequence length (Poli et al., 2023), which allows them to process millions of tokens (Nguyen et al., 2023). However, even in state-space models, very long sequences will impose considerable memory and computational demands.
@@
Bolya et al. (2023), have shown that the efficiency of ViTs can be substantially improved by merging tokens throughout the transformer architecture. Specifically, they compute similarity scores between tokens and combine them into single tokens through a convex combination. However, they only explore token merging for ViT architectures.
@@
In this work, we for the first time explore token merging within the time series domain. We introduce a novel local token merging algorithm whose computational complexity varies from quadratic to linear, based on the neighborhood considered for each token merge. This allows token merging to scale to long sequences and be applicable to state-space models. Further, our local merging preserves causality and is the first viable token merging scheme for transformer decoders. The algorithm is illustrated in figure 1. Through comprehensive empirical evaluations, we analyze the impact of token merging on various time series transformer models and state-space models. Our key contributions are:
Token merging in time series We extend token merging from computer vision to time series analysis and propose local merging as a domain-specific token merging algorithm.
Model acceleration Across five time series transformer architectures (5.1), foundation models (5.3), two state-space models (5.4), and six datasets, token merging reveals substantial computational savings with only slight reductions in accuracy. In some settings, it even improves forecasting performance while accelerating models simultaneously. Token merging enhances model throughput by up to and improves forecasting performance by up to .
@@
Token merging outcomes We identify three distinct outcomes when using token merging: 1) a consistent decline in performance when merging more tokens, 2) initial improvements in accuracy with few merged tokens followed by a drop as merging increases, and 3) scenarios where accuracy remains unchanged regardless of the token merging rate.
Understand token merging Our detailed analysis reveals that token merging acts as an adaptive low-pass filter, selectively reducing noise. We further identify model- and dataset-specific properties explaining the effectiveness of our token merging algorithms.
@@ 2 Related work
Time series transformers Recently, many transformer architectures with inductive biases for time series have been proposed. Most of them reduce complexity by modifying the attention mechanism. Informer uses ProbSparse attention (Zhou et al., 2021), while Autoformer leverages autocorrelation as a sequence-based similarity measure (Wu et al., 2021). FEDformer uses the frequency domain to model time series effectively (Zhou et al., 2022). Non-stationary Transformers mitigate the effect of the time series distribution changing over time (Liu et al., 2022b).
@@
Due to their success in the vision and NLP domain, transformer-based foundation models have lately emerged for time series, often used in zero-shot settings. Many works focus on training transformers directly on large and diverse time series datasets, usually with billions of tokens (Garza & Mergenthaler-Canseco, 2023; Das et al., 2023; Rasul et al., 2023; Woo et al., 2024). Inspired by the success of foundation models in NLP, the recently proposed Chronos model converts continuous time series data into a fixed vocabulary (Ansari et al., 2024).
@@
State-space models Due to the quadratic scaling of the attention mechanism, transformer architectures suffer from significant computational cost when processing long sequences. Recently, state-space models have shown promising results in overcoming this challenge. Linear state-space layers solve the sequential processing requirement of RNNs (Gu et al., 2021). The S4 model reduces memory requirements by conditioning the state-space matrix with a low-rank correction (Gu et al., 2022). By using implicit convolutions and a data-aware gating mechanism, Hyena (Poli et al., 2023) became one of the first state-space model architectures to match transformers on NLP tasks. Later work uses hardware-aware algorithms to improve the performance on modern accelerators (Gu & Dao, 2023).
@@
Reducing tokens Many works reduce the number of processed tokens to increase the efficiency of transformers in computer vision and NLP, often by pruning (Meng et al., 2022; Goyal et al., 2020). Marin et al. (2021) merge tokens in ViT architectures to reduce the loss of information associated with pruning. Bolya et al. (2023) enhance the token merging algorithm, which they successfully apply to already trained encoder-only models. Besides initial work on classification tasks (Bolya et al., 2023), subsequent work applies token merging to diffusion models (Bolya & Hoffman, 2023). Kim et al. (2024) combine merging and pruning, while other works investigate optimal merging and pruning rates (Bonnaerens & Dambre, 2023; Chen et al., 2023). Concurrent work adapts token merging to preserve the spectral properties of the token space (Tran et al., 2024). However, their merging algorithm still has quadratic complexity, making it unsuitable for long sequence processing.
Sparse attention and token skipping Besides reducing the number of tokens, sparse attention (Child et al., 2019) and token skipping (Raposo et al., 2024) also decrease the computational requirements of transformer models. In contrast to token merging, sparse attention can only accelerate the attention mechanism itself and not the subsequent MLP, which can take over of the total computation (Marin et al., 2021). Concurrent work, such as token skipping (Raposo et al., 2024), involves the selection of a subset of tokens to be processed in a transformer layer. However, it has only been shown in NLP when training from scratch. Token merging, however, can accelerate already trained models and does not require any training data or fine-tuning. This is especially important for recent foundation models. In our experiments in sections 5.1 and 5.2, token merging successfully accelerates Informer and Autoformer, which already employ sparse attention. We therefore consider token merging as an orthogonal approach.
Here, we propose the first token merging algorithm for the time series domain, which extends beyond previous investigations in ViTs (Bolya et al., 2023; Bolya & Hoffman, 2023). We systematically evaluate the potential to reduce computational effort in time-series-specific architectures. See appendix A for more details on related work.
@@ 3 Token merging
Despite recent advances in efficient transformers, processing long input sequences still induces considerable memory requirements and computational effort. To address this, we first extend token merging, which successfully boosts throughput in computer vision, to time series models. Next, we propose local merging, a domain-specific and efficient token merging algorithm for state-space models and long sequence processing. Finally, we introduce causal merging as a special case of local merging to allow for token merging in decoder architectures and propose dynamic merging to further improve token merging in real-world settings.
@@
Global token merging for time series Let a neural network consist of layers denoted as , where each layer takes the output of the previous layer as input. We assume that the input consists of tokens with dimension . Thereby, the input tokens are generated by a tokenizer out of -dimensional input data . We assume the input consists of time stamps with variates, and .
@@
To improve the computational efficiency of token-based time series models, we extend global token merging from computer vision to the time series domain. Following Bolya et al. (2023), we combine the most similar tokens in each layer, reducing the tokens to be processed in layer to . For this, we split the set of all tokens into two disjoint subsets in alternation to avoid merging conflicts and allow for a parallelized computation of merging correspondences. Here and contain elements each, denoted as and respectively. We compute the cosine similarity between all tokens in both subsets and merge the most similar correspondences by averaging the tokens accordingly. This results in a global token merging algorithm with quadratic complexity. Lastly, Bolya et al. (2023) use a fixed to enable batch processing without needing to pad individual batch elements to the same shape after token reduction. Later, we introduce more adaptive merging schemes through dynamic merging.
@@
Local token merging for time series In this work, we design new token merging mechanisms for time series architectures and demonstrate run-time and even performance improvements over various datasets and models.
@@
Previous work on token merging in image processing explored global merging schemes, where every token of each subset and could be merged with each other (Bolya et al., 2023; Bolya & Hoffman, 2023).
@@
However, computing the similarity between both sets of tokens has a complexity of , which is suboptimal for sequential data often consisting of long token sequences (Godahewa et al., 2021; Grešová et al., 2023), and state-space models featuring subquadratic complexity (Poli et al., 2023; Nguyen et al., 2023).
@@
Therefore, we propose local merging - a superset of token merging - by introducing as a locality constraint where we compute the similarity only on a local subset of tokens:
Figure 1 illustrates the proposed merging algorithm. The locality constraint reduces the complexity to:
Varying the locality, we achieve linear complexity by considering only neighboring tokens for merging up to quadratic complexity by considering a global merging pool, possibly exploiting more redundancy.
@@
For efficient computation, we refactor into a rectangular tensor. An upper bound for the resulting speed-up can be given by . The acceleration of deeper models is expected to increase as more subsequent layers can profit from already merged tokens.
@@
Some time series transformers use processing mechanisms that require a minimum number of tokens in the forward pass. To universally enable token merging in these architectures, we further introduce as the minimum number of remaining tokens. When encountering odd numbers of tokens , we exclude the most recent token from merging, as we expect it to contain the most relevant information following the Markov assumption.
We derive the complexity of the token merging procedures in section B.1 and further discuss the interplay of time-series-specific inductive biases and token merging in section B.2.
Causal token merging for decoders Existing merging schemes are not suitable for causal operations, as global token merging transfers information over arbitrary ranges.
To remedy this limitation and enable token merging in transformer decoders, such as for recent decoder-only foundation models (Das et al., 2023) and encoder-decoder architectures (Ansari et al., 2024), we propose a special case of local merging: By restricting the merging neighborhood to only adjacent tokens with , local merging preserves temporal causality.
Token merging reduces the number of tokens to be processed throughout the model. However, many architectures require a fixed number of decoder output tokens or fixed dimensions for linear projection output layers. To maintain a constant output dimensionality while merging tokens to speed-up the decoder, we unmerge all tokens in a final step. Coherent with our causal merging operation, we clone a previously merged token into two neighboring identical ones, to unmerge it. Bolya & Hoffman (2023) propose an unmerging algorithm for computer vision. However, they only leverage non-causal global token merging. Moreover, they immediately unmerge after every merge, making it unsuitable for long sequence processing, as it is unable to utilize the cumulative effect of reducing tokens.
Dynamic token merging A fixed merging scheme allows for batch processing without needing to pad individual time series to the same length. However, it enforces a constant among layers and batches, which might not always be optimal, as dissimilar tokens might be merged. We introduce dynamic merging to further improve token merging in real-world settings, where batch sizes are often small. To this end, we determine the number of tokens to be merged dynamically for every batch element using a token cosine similarity threshold. To avoid padding, we average them throughout the batch. As a result, dynamic merging enables optimal merging rates that are adaptive to the input data and current network layer.
@@ 4 Experiments
We systematically explore local merging in diverse settings on time series datasets and model architectures in different sizes each. Additionally, we investigate local merging in large foundation models using Chronos in a zero-shot setting (Ansari et al., 2024). Finally, we demonstrate that local merging can be applied to state-space models for long sequence processing. See appendix C for more details on our experimental settings.
Datasets We use time series forecasting datasets including ETTh1, ETTm1, Weather, Electricity and Traffic for our transformer experiments. For state-space models, we use the long-range Dummy Mouse Enhancers Ensembl dataset.
Model architectures For our main experiments, we use architectures, including Autoformer, FEDformer, Informer, Non-stationary Transformer, and the vanilla Transformer (Vaswani et al., 2017) as reference. For each model, we evaluate local merging for different model sizes with encoder layers, which we train doing hyperparameter optimization. We use an input length of , following the results of Nie et al. (2023), and a prediction horizon samples. Longer sequences would generally benefit token merging.
For experiments on the foundation model Chronos, we use the default input length of and prediction horizon (Ansari et al., 2024). We compute the median from Chronos probabilistic forecasts and report the MSE.
For our experiments on state-space models, we use HyenaDNA medium, a genomic foundation model (Nguyen et al., 2023) based on the Hyena architecture (Poli et al., 2023) and Mamba (Gu & Dao, 2023) models with the same hyperparameters as Hyena. We use a large input length of nucleotides, utilizing state-space models’ subquadratic complexity.
Applying local merging Allowing self-attention to transfer information between tokens before merging them is beneficial in our experiments. Therefore, we apply local merging after the self-attention in transformer architectures as Bolya et al. (2023). In state-space models, we merge tokens after the Hyena or Mamba operator.
Reproducibility of measurements We report all results on the same Nvidia A6000 GPU and do multiple measurements to achieve inference time standard deviations .
@@ 5 Results
We first present our main results for local merging on pretrained models and models trained with local merging. Next, we scale local merging to large foundation models and explore token merging for state-space models and long sequence processing. Finally, we investigate if we can gain even higher speed-ups in real-world settings, leveraging dynamic merging schemes.
5.1 Local merging in pretrained models
We investigate local merging in both the encoder and decoder on diverse time series transformer models with different inductive biases. All models are trained on the target dataset and local merging is applied only during inference time, as accelerating already trained models is of high practical relevance. We choose local merging hyperparameters as described in appendix C, selecting the fastest token merging trial on the validation set that is within a increase in MSE compared to the reference without token merging.
If we do not find a trial with token merging satisfying these tight criteria, we report results without token merging, mimicking how local merging might be applied in practice. We perform all selections on the validation set and report all results on the test set.